xi 1 2 3 4 5
yi 4 7 6 12 15

d. Develop the estimated regression equation by computing the values of b0 an b1
y=_____ +______x
e. Use of estimated regression equation to predict the value of y when x =5
y=_____
Answer #1

NE hi 지 4 나 2. 구 나. 14 3 1 4 18 나 12. IC 48 5 5 25 15 내 55. 159 b0 = 파이지 bl = Cov(9) var (2) il AI 금차 Hect, X-1-3, 3- - S8, 는so, the regression equation is

y=0.7+(2.7 * x)

e) predicted value of y when x=5 is

y= 0.7+(2.7*5) = 14.2.